D14610:

The table describes the quadratic function h(x). x h(x) −3 1 −2 −2 −1 −3 0 −2 1 1 2 6 3 13 What is the equation of h(x) in vertex form?

toga:

To find the equation of h(x) in vertex form, we need to complete the square. First, we can find the axis of symmetry by averaging the x-values of the two points on either side of the vertex: Axis of symmetry = (−2 + (−1))/2 = −1.5 This tells us that the x-coordinate of the vertex is −1.5. To find the y-coordinate, we can substitute this value into the function: h(−1.5) = −3 So the vertex is at (−1.5, −3). Now we can write the equation in vertex form: h(x) = a(x − (−1.5))2 − 3 To find the value of a, we can use one of the other points on the graph. Let's use (2, 6): 6 = a(2 − (−1.5))2 − 3 9 = 5.25a a = 9/5.25 = 1.71 (rounded to two decimal places) So the equation of h(x) in vertex form is: h(x) = 1.71(x + 1.5)2 − 3 I hope this helps!

surjithayer:

let \[h(x)=ax^2+bx+c~~ ...(1)\] when x=0, h(x)=-2 put in (1) \[-2=a(0)^2+b(0)+c\] c=-2 \[h(x)=ax^2+bx-2~~...(2)\] when x=-1,h(x)=-3 \[-3=a(-1)^2+b(-1)-2\] -3=a-b-2 a-b=-3+2=-1 a-b=-1 ...(3) when x=1,h(x)=1 \[1=a(1)^2+b(1)-2\] a+b=1+2 a+b=3 ...(4) add (3) and (4) 2a=-1+3 2a=2 a=1 from (4) 1+b=3 b=3-1=2 so function is \[h(x)=x^2+2x-2\] \[h(x)=x^2+2x+1-1-2\] \[h(x)=(x+1)^2-3\] which is in vertex form and vertex is (-1,-3)

Prettygirl099:

jeez

LILBOY:

idk

toga:

@lilboy wrote:
idk
@lilboy wrote:
idk
this is spam